OpenStudy (anonymous):

(-1,9/5) and (7/3,2/5)

OpenStudy (anonymous):

@Hero

OpenStudy (anonymous):

A. (4/3,11/5) B. (-5/3,7/10) C. (5/3,-7/10) D. (2/3,11/10)

OpenStudy (anonymous):

Find the midpoint of the line segment whose end points are given.

OpenStudy (anonymous):

midpoint of (a,b) (c,d) is \[(\frac{ a+c }{ 2 }, \frac{ c+d }{ 2 })\]

OpenStudy (anonymous):

HUH?

OpenStudy (anonymous):

You add the two X-values and divide it by two, that's the middle of the x coordinates. Now do the same for Y. Add the two Y values and then divide by two.

OpenStudy (anonymous):

The coordinate you get from this would be the "mid-point"

OpenStudy (anonymous):

how do i do it with fractions in the way??????

OpenStudy (anonymous):

@Hero

OpenStudy (anonymous):

\[\frac{ -6 }{ 3 }\div2 ,\frac{ 11 }{ 5 }\div2\]

OpenStudy (anonymous):

 This is how you divide with fractions. Look at it this way then multiply.